# Slot Attention with Re-Initialization and Self-Distillation

### v1

@@ 3.3. Self-Distilled Aggregation
The key to realizing self-distillation or simply distillation is that the teacher should be better than the student during training (of course after training the student can be better than the teacher). SPOT (Kakogeorgiou et al., 2024) resorts to pretrain an extra copy of the model as the teacher to realize such performance gap.
@@
SPOT (Kakogeorgiou et al., 2024) utilizes all the decoder attention map as the target without selection. And if some decoding attention compositions are worse than the corresponding aggregation attention, then such approximation would harm the performance. This is also why they have to pretrain an extra copy of the model and use it as the teacher to guide the training of the (student) model. Such offline distillation at least doubles the training cost.
@@ 3.4. Random Auto-Regressive Decoding
As discussed in Section 2 “OCL decoding”, three types OCL decoding all have their disadvantages. Mixture-based decoders like CNN (Kipf et al., 2022; Elsayed et al., 2022) and MLP (Seitzer et al., 2023) are too computationally expensive, while SlotMixer (Zadaianchuk et al., 2024) is efficient but has limited object representation quality. Diffusion-based decoders (Jiang et al., 2023; Wu et al., 2023b) are expensive and sensitive to train. Auto-regression-based decoders like Transformer decoder (Singh et al., 2022a, b) are very efficient, but their next-token prediction has to flatten the 2-dimensional super-pixels into 1-dimensional in a fixed order, thus they do not utilize the spatial correlation among super-pixels. Transformer9 (Kakogeorgiou et al., 2024) utilizes nine handcrafted permutations $\{\bm{I}^{j}|j=1...9\}$ to better capture spatial dependencies:
@@
where $\bm{I}^{j}\in\mathbb{R}^{n}$ is the $j$th handicraft super-pixel ordering, with $n=h\times w$; $\bm{E}^{j}_{\mathrm{bos}}\in\mathbb{R}^{c}$ is the $j$th Beginning-of-Sentence (BoS) token (Kakogeorgiou et al., 2024); $\mathrm{flatten}_{h,w}(\cdot)$ flattens height and width; $\mathrm{gather}_{n}(\cdot,\cdot)$ select the tensor along dimension $n$ given index tensor $\bm{I}^{j}$; $\mathrm{concat}_{n}(\cdot,\cdot)$ concatenates two tensors along dimension $n$; $\bm{Y}^{j}_{\mathrm{bos}}\in\mathbb{R}^{n\times c}$ is the sequence prepended with BoS; $\bm{M}_{\mathrm{causal}}\in\mathbb{R}^{n\times n}$ is the typical causal mask (Kakogeorgiou et al., 2024); and $\bm{\phi}_{\mathrm{d}}$ is parameterized as a Transformer decoder.
@@ 4.1. Object Discovery
The attention maps of OCL aggregation and decoding are the segmentation of the objects and background that are discovered by slots. We measure these segmentation using typical metrics, i.e., ARI (Adjusted Rand Index) (Learn, 2025a), ARIfg (foreground Adjusted Rand Index), mBO (mean Best Overlap) (Uijlings et al., 2013) and mIoU (mean Intersection-over-Union) (Learn, [n. d.]). ARI measures more about larger areas, i.e., the background; mBO only measures the best overlapped segmentation; mIoU is the most strict measurement.
For baselines, we use the models from (Zhao et al., 2025b), which reproduce very strong baselines by unifying representative OCL models all with current best practices. Among them, SLATE, STEVE, DINOSAUR, SlotDiffusion and SPOT are designed for images; STEVE and VideoSAUR are for videos; VVO with different decoders support images or videos. We also include AdaSlot (for images) (Fan et al., 2024) and SOLV (for videos) (Aydemir et al., 2023).
For datasets, we use ClevrTex (the main set for training and the out-of-distribution set for evaluation) (Karazija et al., [n. d.]), COCO (instance segmentation) (Lin et al., 2014), VOC (Everingham et al., 2010), MOVi-D (Research, [n. d.]) and YTVIS (high-quality) (SysCV, [n. d.]). Among them, ClevrTex is synthetic images with complex textures; COCO and VOC are real-world images; MOVi-D is a set of challenging synthetic videos with up to 20 objects per scene; YTVIS is real-world YouTube videos.
As shown in Table 1, with unconditional query, our DIAS outperforms all the baselines on both synthetic and real-world images in object discovery under all the metrics. In fact, DIAS achieves new state-of-the-art performance. As shown in Table 2, with conditional query, DIAS the temporal version surpasses STEVE by a large margin on synthetic videos; With unconditional query, DIAS defeats the state-of-the-art video OCL model. As shown in Table 3, combined with the general improver VVO (Zhao et al., 2025b), DIAS pushes state-of-the-art performance forward even further.
@@
Table 2. Object discovery on videos. Using DINO2 ViT s/14 for OCL encoding; input resolution is 256$\times$256 (224$\times$224).
@@
Table 3. Combined with technique VVO, DIAS pushes SotA forward even further. Using DINO2 ViT s/14 for OCL encoding; input resolution is 256$\times$256 (224$\times$224).
@@ 4.2. Object Recognition
With slots representing different objects or background, we can further evaluate how much object information they aggregate. In other studies, this is also called set prediction (Locatello et al., 2020) or object property prediction (Fan et al., 2024). Following the routine of (Seitzer et al., 2023), we reuse the trained models from Section 4.1 to extract the real-world image dataset COCO into slots representation, and train a small MLP to take each slot as input to predict the corresponding object’s class label and bounding box. We use metrics top1 (Learn, 2025d) and R2 (Learn, 2025c) to measure the classification and regression performance.
@@ 4.3. Visual Prediction and Reasoning
With improved object representation, our DIAS also helps the downstream tasks, like visual prediction and reasoning. Following the rountine from (Wu et al., 2023a), we firstly train our DIAS and other baselines on Physion (Bear et al., 2021), which is a synthetic video dynamics modeling dataset, and represent the dataset as slots. Then we train an object-centric world model SlotFormer (Wu et al., 2023a) to auto-regressively predict future frames of slots, given initial five frames of slots as the input. And based on such predictions, we further train a reasoning model to classify whether the stimuli objects will contact finally.

### v3

@@ 3.3. Self-Distilled Aggregation
The key to realizing self-distillation or simply distillation is that the teacher should be better than the student during training (of course after training the student can be better than the teacher). SPOT (kakogeorgiou2024spot) resorts to pretrain an extra copy of the model as the teacher to realize such performance gap.
@@
SPOT (kakogeorgiou2024spot) utilizes all the decoder attention map as the target without selection. And if some decoding attention compositions are worse than the corresponding aggregation attention, then such approximation would harm the performance. This is also why they have to pretrain an extra copy of the model and use it as the teacher to guide the training of the (student) model. Such offline distillation at least doubles the training cost.
@@ 3.4. Random Auto-Regressive Decoding
As discussed in Section 2 “OCL decoding”, three types OCL decoding all have their disadvantages. Mixture-based decoders like CNN (kipf2021savi; elsayed2022savipp) and MLP (seitzer2023dinosaur) are too computationally expensive, while SlotMixer (zadaianchuk2024videosaur) is efficient but has limited object representation quality. Diffusion-based decoders (jiang2023lsd; wu2023slotdiffuz) are expensive and sensitive to train. Auto-regression-based decoders like Transformer decoder (singh2021slate; singh2022steve) are very efficient, but their next-token prediction has to flatten the 2-dimensional super-pixels into 1-dimensional in a fixed order, thus they do not utilize the spatial correlation among super-pixels. Transformer9 (kakogeorgiou2024spot) utilizes nine handcrafted permutations $\{\bm{I}^{j}|j=1...9\}$ to better capture spatial dependencies:
@@
where $\bm{I}^{j}\in\mathbb{R}^{n}$ is the $j$th handicraft super-pixel ordering, with $n=h\times w$; $\bm{E}^{j}_{\mathrm{bos}}\in\mathbb{R}^{c}$ is the $j$th Beginning-of-Sentence (BoS) token (kakogeorgiou2024spot); $\mathrm{flatten}_{h,w}(\cdot)$ flattens height and width; $\mathrm{gather}_{n}(\cdot,\cdot)$ select the tensor along dimension $n$ given index tensor $\bm{I}^{j}$; $\mathrm{concat}_{n}(\cdot,\cdot)$ concatenates two tensors along dimension $n$; $\bm{Y}^{j}_{\mathrm{bos}}\in\mathbb{R}^{n\times c}$ is the sequence prepended with BoS; $\bm{M}_{\mathrm{causal}}\in\mathbb{R}^{n\times n}$ is the typical causal mask (kakogeorgiou2024spot); and $\bm{\phi}_{\mathrm{d}}$ is parameterized as a Transformer decoder.
@@ 4.1. Object Discovery
The attention maps of OCL aggregation and decoding are the segmentation of the objects and background that are discovered by slots. We measure these segmentation using typical metrics, i.e., ARI (Adjusted Rand Index) (sklearn2025ari), ARIfg (foreground Adjusted Rand Index), mBO (mean Best Overlap) (uijlings2013selectivesearch) and mIoU (mean Intersection-over-Union) (sklearn2025miou). ARI measures more about larger areas, i.e., the background; mBO only measures the best overlapped segmentation; mIoU is the most strict measurement.
For baselines, we use the models from (zhao2025vvo), which reproduce very strong baselines by unifying representative OCL models all with current best practices. Among them, SLATE, STEVE, DINOSAUR, SlotDiffusion and SPOT are designed for images; STEVE and VideoSAUR are for videos; VVO with different decoders support images or videos. We also include AdaSlot (for images) (fan2024adaslot) and SOLV (for videos) (aydemir2023solv).
For datasets, we use ClevrTex (the main set for training and the out-of-distribution set for evaluation) (karazija2clevrtex), COCO (instance segmentation) (lin2014coco), VOC (everingham2010voc), MOVi-D (google2025movi) and YTVIS (high-quality) (syscv2022hqytvis). Among them, ClevrTex is synthetic images with complex textures; COCO and VOC are real-world images; MOVi-D is a set of challenging synthetic videos with up to 20 objects per scene; YTVIS is real-world YouTube videos.
As shown in Table 1, with unconditional query, our DIAS outperforms all the baselines on both synthetic and real-world images in object discovery under all the metrics. In fact, DIAS achieves new state-of-the-art performance. As shown in Table 2, with conditional query, DIAS the temporal version surpasses STEVE by a large margin on synthetic videos; With unconditional query, DIAS defeats the state-of-the-art video OCL model. As shown in Table 3, combined with the general improver VVO (zhao2025vvo), DIAS pushes state-of-the-art performance forward even further.
@@
Table 2. Object discovery on videos. Using DINO2 ViT s/14 for OCL encoding; input resolution is $\times256$ ($\times224$).
@@
Table 3. Combined with technique VVO, DIAS pushes SotA forward even further. Using DINO2 ViT s/14 for OCL encoding; input resolution is $\times256$ ($\times224$).
@@ 4.2. Object Recognition
With slots representing different objects or background, we can further evaluate how much object information they aggregate. In other studies, this is also called set prediction (locatello2020slotattent) or object property prediction (fan2024adaslot). Following the routine of (seitzer2023dinosaur), we reuse the trained models from Section 4.1 to extract the real-world image dataset COCO into slots representation, and train a small MLP to take each slot as input to predict the corresponding object’s class label and bounding box. We use metrics top1 (sklearn2025topkacc) and R2 (sklearn2025r2) to measure the classification and regression performance.
@@ 4.3. Visual Prediction and Reasoning
With improved object representation, our DIAS also helps the downstream tasks, like visual prediction and reasoning. Following the rountine from (wu2022slotformer), we firstly train our DIAS and other baselines on Physion (bear2021physion), which is a synthetic video dynamics modeling dataset, and represent the dataset as slots. Then we train an object-centric world model SlotFormer (wu2022slotformer) to auto-regressively predict future frames of slots, given initial five frames of slots as the input. And based on such predictions, we further train a reasoning model to classify whether the stimuli objects will contact finally.